[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Mon Apr 10, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Mon Apr 10 13:15:24 UTC 2017


HSU Shift Log for Mon Apr 10 2017 09:15:15 GMT-0400 (EDT)

Shift: H

Forecaster: E Blake

Technical Issues Summary:

ATCF issues:

N-AWIPS issues:

AWIPS-2/NCP issues:

Other Techincal issues: Non-operational of course but the METSAT-7 data  
(Indian Ocean) for both the IR/VIS channels are blank and have been for  
about 5 days. I think MetSat 8 is the operational data now (?)

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170410/f1befe1b/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list